# Are We Overlooking the Dimensions? Learning Latent Hierarchical Channel Structure for High-Dimensional Time Series Forecasting

### v1

@@ Abstract
Time series forecasting (TSF) is a central problem in time series analysis. However, as the number of channels in time series datasets scales to the thousands or more, a scenario we define as High-Dimensional Time Series Forecasting (HDTSF), it introduces significant new modeling challenges that are often not the primary focus of traditional TSF research. HDTSF is challenging because the channel correlation often forms complex and hierarchical patterns. Existing TSF models either ignore these interactions or fail to scale as dimensionality grows. To address this issue, we propose U-Cast, a channel-dependent forecasting architecture that learns latent hierarchical channel structures with an innovative query-based attention. To disentangle highly correlated channel representation, U-Cast adds a full-rank regularization during training. We also release Time-HD, a benchmark of large, diverse, high-dimensional datasets. Our theory shows that exploiting cross-channel information lowers forecasting risk, and experiments on Time-HD demonstrate that U-Cast surpasses strong baselines in both accuracy and efficiency. Together, U-Cast and Time-HD provide a solid basis for future HDTSF research. To promote reproducibility and future extensions, we release Time-HD-lib at GitHub, and Time-HD can be accessed at Hugging Face.
@@ 1 Introduction
Figure 1: Time-HD provides diverse high-dimensional datasets.
@@
Time series data involving multiple variables (a.k.a. channels) that change over time are fundamental to numerous real-world applications, such as stock market prices granger2014forecasting , highway traffic flows yin2021deep , solar power generation outputs wan2015photovoltaic , and temperature measurements wu2023interpretable . Consequently, time series forecasting (TSF) has become a central challenge in the time series analysis community, aiming to predict future values based on historical observations chatfield2000time ; de200625 . Recent advances in machine learning have catalyzed significant methodological progress in this field wang2024deep ; lim2021time , leading to a variety of novel architectures autoformer ; fedformer ; informer ; itransformer ; nbeats ; patchtst ; timemixer ; timesnet ; timedistill ; feng2025seizureformer ; zhao2025tat . These models can be broadly categorized by how they handle the inter-channel dependencies: channel-dependent (CD) methods itransformer ; crossformer ; autoformer ; informer ; ccm ; fedformer explicitly capture cross-variable interactions, while channel-independent (CI) methods patchtst ; dlinear ; pdf ; timemixer ; fits ; sparsetsf treat each variable separately.
@@
However, as the number of channels in time series datasets scales into the thousands or more, a scenario we define as High-Dimensional Time Series Forecasting (HDTSF) has posed significant new modeling challenges that are often not the primary focus of traditional TSF research. In the HDTSF regime, the sheer volume of channels can complicate the channel dependency patterns and render traditional models, which are often developed and evaluated on low-dimensional datasets, computationally infeasible or unreliable. More importantly, high dimensionality often gives rise to latent hierarchical structures within the channels. These structures are common in large real-world systems, where channels exhibit implicit groupings and multi-level correlations based on factors such as spatial proximity (e.g., nested geographical regions in climate data) or semantic relationships (e.g., stocks within related economic sub-sectors). While some explicit hierarchies might be known, many are latent and can evolve dynamically. Existing TSF models are often not designed to discover or leverage these complex, multi-scale hierarchical channel dependencies, which may significantly limit their generalizability in high-dimensional contexts.
To address the specific challenge of effectively modeling these latent hierarchical channel structures within HDTSF, we propose U-Cast, a novel and efficient forecasting model that incorporates an innovative query-based attention mechanism within a Hierarchical Latent Query Network and a Hierarchical Upsampling Network. This architecture is expressly designed to infer latent hierarchical organizations among channels and efficiently compute multi-scale correlations. To further enhance the distinctiveness of learned representations within this hierarchy, especially when dealing with highly correlated raw channel data, U-Cast employs a full-rank regularization technique.
Developing and rigorously evaluating methods like U-Cast, which are tailored for the intricacies of HDTSF and its characteristic hierarchical dependencies, requires datasets that adequately reflect these complexities. However, commonly used TSF benchmarks (e.g., ETT informer , Weather autoformer ) often feature a limited number of channels and may not possess the scale or structural richness representative of many emerging high-dimensional applications. This gap can hinder the robust assessment of model scalability, the nuanced benefits of sophisticated dependency modeling strategies, and overall generalizability. To facilitate rigorous research in HDTSF and provide a suitable testbed for U-Cast and future innovative approaches, we have curated Time-HD. Time-HD is the first comprehensive benchmark suite specifically focused on high-dimensional time series, offering large-scale, domain-diverse datasets, as illustrated in Figure 1. Our experiments on Time-HD validate U-Cast’s efficacy in learning hierarchical structures and achieving superior forecasting accuracy in these demanding high-dimensional settings. Our main contributions are summarized as:
C1: Detailed analysis of channel dependency. We provide rigorous theoretical and empirical evidence (Section 3) that highlights the importance of effectively modeling inter-channel dependencies for time series forecasting. It shows that CD can achieve lower Bayes risk than CI, and that incorporating more relevant channels can further improve forecasting accuracy, highlighting the need for high-dimensional benchmark datasets.
@@
C2: A strong model for high-dimensional time series forecasting. We propose U-Cast (Section 4), a new CD forecasting model that employs an innovative query-based attention mechanism to efficiently learn latent hierarchical channel structures and enables scalable modeling of inter-channel correlations. U-Cast combines strong performance across Time-HD with the best efficiency among baselines, making it a reference point for future models.
C3: A large-scale high-dimensional time series forecasting benchmark. We release Time-HD (Section 5), the first comprehensive benchmark suite specifically focused on high-dimensional time series forecasting. These datasets, characterized by substantial scale and broad domain representation, are systematically collected, curated, and preprocessed to offer a realistic and challenging testbed for evaluating model scalability and generalizability. We observe that CD methods do not consistently outperform CI methods, likely because current CD baselines either fail to capture complex channel dependencies or do not scale to high-dimensional time series.
C4: An open-source toolkit for reproducible research. To cultivate a collaborative research ecosystem, we release Time-HD-lib. This open-source library provides an end-to-end framework supporting the Time-HD benchmark, including standardized data preprocessing, seamless integration of datasets, unified evaluation protocols, and facilities for hyperparameter optimization, thus promoting reproducibility and future extensions.
@@ 2 Related Work
High-Dimensional Time Series Datasets. Some high-dimensional datasets have already been used in recent work on building time series foundation models moirai ; chronos ; time-moe ; moment . These studies primarily focus on leveraging high-dimensional datasets for pretraining. However, several limitations exist in how they handle such datasets. First, during pretraining, channels are treated independently, capturing only temporal dependencies without modeling inter-channel correlations. Second, despite pretraining on high-dimensional datasets, fine-tuning and evaluation are still performed on low-dimensional ones. Third, some pretraining datasets have unaligned timestamps, complicating their use in evaluation. These limitations highlight the importance of developing high-dimensional time series datasets specifically tailored for benchmarking and evaluation in time series forecasting research. Table 5 in Appendix A provides a more detailed summarization of the differences between Time-HD and existing works.
Multivariate Time Series Forecasting. Previous methods for modeling multivariate time series can be broadly categorized into channel-independent (CI) and channel-dependent (CD) strategies. The CI modeling strategy employs a shared model backbone across all channels and processes each channel independently during the forward pass patchtst ; dlinear ; fits ; timemixer ; pattn ; pdf ; sparsetsf . This design typically results in lower model capacity but offers greater robustness. In contrast, the CD modeling strategy introduces dedicated modules to capture inter-channel dependencies. Based on the granularity at which these correlations are modeled, CD can be further divided into position-wisetsmixer ; timesnet ; autoformer ; informer ; fedformer ; crossformer , token-wiseitransformer ; timexer , and cluster-wise ccm ; duet approaches. These methods are expected to provide higher capacity and make better use of cross-channel information. However, due to the constraints of existing low-dimensional multivariate time series forecasting benchmarks, most methods have not been validated at scale. We provide a more detailed discussion in Appendix A.
@@ 3 Preliminary Study
Prior studies on existing time series benchmarks, which are often of low dimensionality, do not consistently show a clear advantage of CD over CI models zhao2024rethinking ; patchtst ; dlinear ; pdf ; zhao2025timerecipe . This observation might lead to skepticism regarding the practical benefits of explicitly modeling inter-channel correlations. However, we hypothesize that the limited dimensionality of these benchmarks inherently restricts the potential gains from CD approaches. This preliminary study aims to theoretically and empirically investigate the impact of data dimensionality on the forecasting performance achievable by leveraging channel correlations. We begin with a theoretical analysis of risk reduction.
@@ 3.1 Theoretical Analysis of Risk Reduction from CD Models
To formally analyze the benefits of CD modeling, we consider a simplified setting. Assume a bivariate time series ($C=2$) generated by a Vector Autoregression process of order 1 (VAR(1)) stock2001vector :
@@
For the time series described in Eq. (1), the Bayes risks of CI and CD models under squared-error loss satisfy: $R_{\text{CI}}-R_{\text{CD}}\geq 0$. This inequality is strict if and only if $a_{12}\neq 0$ and $a_{21}\neq 0$, and the conditional variances $\operatorname{Var}(z_{t}^{(1)}\mid z_{t}^{(2)})$ (representing the variance of $z_{t}^{(1)}$ that remains after $z_{t}^{(2)}$ is known) and $\operatorname{Var}(z_{t}^{(2)}\mid z_{t}^{(1)})$ are positive.
We provide the proof of this theorem in Appendix E. Theorem 1 establishes that CD modeling is at least as good as CI modeling. A strict advantage for CD arises if: (1) there is genuine predictive information flowing between channels (non-zero $a_{12}$ or $a_{21}$ means one channel directly influences the future of the other), and (2) each channel contains some unique information not present in the other (positive conditional variances). If channels are truly independent in their evolution ($a_{12}=a_{21}=0$) or one is a deterministic function of the other, then $R_{\text{CI}}=R_{\text{CD}}$.
To investigate the impact of increasing dimensionality, we now consider a general $P$-channel VAR(1) process: $z_{t+1}=\mathbf{A}z_{t}+\varepsilon_{t+1}$, where $z_{t}\in\mathbb{R}^{P}$. We focus on forecasting a single target channel, say $Y=z_{t+1}^{(1)}$, which is given by $z_{t+1}^{(1)}=\sum_{j=1}^{P}a_{1j}z_{t}^{(j)}+\varepsilon_{t+1}^{(1)}$, with $\operatorname{Var}(\varepsilon_{t+1}^{(1)})=\sigma_{11}$. Let $R_{p}$ denote the Bayes risk (minimum MSE) for forecasting $Y=z_{t+1}^{(1)}$ using the information from the first $p$ channels at time $t$, i.e., $\{z_{t}^{(1)},\dots,z_{t}^{(p)}\}$.
@@
The proof is provided in Appendix F. Theorem 2 formally shows that incorporating additional informative and non-redundant channels progressively reduces the forecasting risk for a target variable. This underscores the theoretical benefit of using more relevant dimensions (channels), supporting our hypothesis that higher-dimensional data provides greater opportunity for CD models to outperform CI approaches.
@@ 3.2 Empirical Analysis
Data Generation. We generate synthetic multivariate time series using a $C$-dimensional Vector Autoregression process in Equation 1, where the coefficient matrix ${\bf A}\in\mathbb{R}^{C\times C}$ encodes the dependency structure of channels. We focus on two distinct structures for ${\bf A}$: (1) Independent: Channels evolve independently of each other i.e., ${\bf A}$ is a diagonal matrix. For this setting, we use $C=100$ channels. (2) Anti-Self: Channels depend significantly on other channels but not on their own immediate past, i.e., ${\bf A}$ has zero diagonal entries and non-zero off-diagonal entries. To examine the impact of dimensionality, we generate Anti-Self datasets with $C\in\{100,250,2000\}$. More experimental details are provided in Appendix I.
@@
Results and Discussion. As shown in Table 1, the CI model performs better in the Independent setting, where no inter-channel dependencies exist. In contrast, in all Anti-Self settings, the CD model achieves substantially lower MSE with larger gains as the number of interacting channels increases. These empirical findings are consistent with our theoretical analysis: CD modeling becomes increasingly effective when significant inter-channel dependencies exist, especially in higher-dimensional systems. These observations highlight a pressing need for progress on two critical fronts. First, there is a clear demand for novel CD models capable of efficiently and effectively navigating the intricacies of high-dimensional time series data. Second, the development of dedicated high-dimensional forecasting benchmarks, characterized by complex channel dependencies, is crucial for the rigorous evaluation and continued advancement of such sophisticated methods. Accordingly, in the subsequent sections, we address these needs by first introducing our U-Cast framework specifically tailored for HDTSF and then presenting Time-HD, a comprehensive benchmark designed to support and galvanize further research in this evolving domain.

### v2

# U-Cast: Learning Hierarchical Structures for High-Dimensional Time Series Forecasting

@@ Abstract
Time series forecasting (TSF) is a central problem in time series analysis. However, as the number of channels in time series datasets scales to the thousands or more, a scenario we define as High-Dimensional Time Series Forecasting (HDTSF), it introduces significant new modeling challenges that are often not the primary focus of traditional TSF research. HDTSF is challenging because the channel correlation often forms complex and hierarchical patterns. Existing TSF models either ignore these interactions or fail to scale as dimensionality grows. To address this issue, we propose U-Cast, a channel-dependent forecasting architecture that learns latent hierarchical channel structures with an innovative query-based attention. To disentangle highly correlated channel representation, U-Cast adds a full-rank regularization during training. We also release Time-HD, the first benchmark of large, diverse, high-dimensional datasets. Our theory shows that exploiting cross-channel information lowers forecasting risk, and experiments on Time-HD demonstrate that U-Cast surpasses strong baselines in both accuracy and efficiency. Together, U-Cast and Time-HD provide a solid basis for future HDTSF research. Our code and benchmark are available to ensure reproducibility.
@@ 1 Introduction
Table 1: Comparison of existing datasets by channel scale.
Datasets #Channels
ETT 7
Weather 21
Solar 137
ECL 321
Traffic 862
Time-HD 1k–20k
@@
Time series data involving multiple variables (a.k.a. channels) that change over time are fundamental to numerous real-world applications. Consequently, time series forecasting (TSF) has become a central challenge in the time series analysis community, aiming to predict future values based on historical observations (Chatfield, 2000; De Gooijer & Hyndman, 2006). Recent advances in machine learning have catalyzed significant methodological progress in this field (Wang et al., 2024b; Lim & Zohren, 2021), leading to a variety of novel architectures (Wu et al., 2021; Zhou et al., 2022; 2021; Liu et al., 2024a; Oreshkin et al., 2020; Nie et al., 2023; Wang et al., 2024a; Wu et al., 2022; Ni et al., 2025).
Despite this progress, most existing TSF datasets, as shown in Table 1 and Table 6, remain relatively low-dimensional, with only a handful of channels (e.g., ETT, Weather) to at most a few hundred (e.g., Traffic, ECL). However, in many real-world applications, the number of channels can easily scale into the thousands or more. For instance, financial markets involve thousands of stocks (Granger & Newbold, 2014), urban traffic systems rely on tens of thousands of sensors (Yin et al., 2021), smart grids generate massive streams from large-scale meters (Wan et al., 2015), and climate reanalysis contains hundreds of thousands of variables defined on global spatial grids (Wu et al., 2023). We define this scenario as High-Dimensional Time Series Forecasting (HDTSF), where the scale of channels introduces significant new modeling challenges that are often not the primary focus of traditional TSF research.
@@
Two difficulties are especially pressing. First, scaling to thousands of channels with complicated dependencies makes many channel-dependent models, such as iTransformer (Liu et al., 2024a) and TSMixer (Chen et al., 2023), computationally inefficient or unreliable. Second, high dimensionality often gives rise to latent hierarchical structures among the channels. These structures are common in large real-world systems (Appendix M), where channels exhibit implicit groupings and multi-level correlations based on factors such as spatial proximity (e.g., nested geographical regions in climate data) or semantic relationships (e.g., stocks within related economic sub-sectors). While such multi-scale correlations are pervasive, most existing models are not designed to explicitly discover or leverage them, which limits generalizability in high-dimensional contexts. This motivates two central research questions: (1) How can we build forecasting architectures that explicitly capture latent hierarchical structures to scale across thousands of channels? (2) How can we systematically evaluate such models when existing benchmarks rarely exceed a few hundred channels?
To address the first question, we propose U-Cast, a scalable forecasting architecture that introduces a Hierarchical Latent Query Network and a Hierarchical Upsampling Network to efficiently uncover latent multi-scale channel organizations. To disentangle highly correlated inputs, U-Cast also employs a full-rank regularization objective that encourages diverse, non-redundant channel representations. Answering the second question requires a new empirical foundation. Existing TSF benchmarks contain only limited channels and may not possess the scale or structural richness representative of high-dimensional applications, making them inadequate for evaluating scalability and advanced dependency modeling. To this end, we curate Time-HD, the first comprehensive benchmark suite for HDTSF. Time-HD spans 16 datasets with 1k–20k channels across diverse domains, providing a much-needed testbed for evaluating scalability and dependency modeling. While this paper does not claim to exhaustively analyze all benchmarking results, we release Time-HD as a community resource to facilitate systematic study of HDTSF in future work. Our experiments on Time-HD validate U-Cast’s efficacy in learning hierarchical structures and achieving superior forecasting accuracy in these demanding high-dimensional settings. Our main contributions are summarized as:
C1: Theoretical and empirical analysis of high-dimensional channel dependency. We prove that channel-dependent (CD) models have lower Bayes risk than channel-independent (CI) models whenever non-redundant channels exist, with the advantage increasing in higher dimensions (Section 3). Our controlled experiments on synthetic and real-world datasets further corroborate these findings.
@@
C2: U-Cast: A scalable architecture for HDTSF. We propose U-Cast (Section 4), a new CD forecasting model that introduces an innovative query-based attention mechanism to efficiently learn latent hierarchical channel structures and enables scalable modeling of inter-channel correlations. U-Cast combines strong performance across Time-HD with the best efficiency among baselines, making it a strong reference design for future HDTSF models.
C3: Time-HD: the first benchmark suite for HDTSF. We release Time-HD (Section 5), the first comprehensive benchmark suite curated for HDTSF. Time-HD spans 16 datasets (1k–20k channels) across diverse domains, providing the necessary foundation for evaluating scalability and dependency modeling in realistic high-dimensional contexts. While not intended as an exhaustive benchmarking study, Time-HD establishes a standardized and challenging testbed that enables rigorous evaluation and comparison of future models.
C4: Open-source library for reproducibility. We provide Time-HD-lib, an open-source library supporting the benchmark. It offers standardized preprocessing, unified evaluation protocols, and automated hyperparameter optimization. This toolkit ensures reproducibility of our results and lowers the barrier for future work on HDTSF.
@@ 2 Related Work
High-Dimensional Time Series Datasets. Some high-dimensional datasets have already been used in recent work on building time series foundation models (Woo et al., 2024; Ansari et al., 2024; Shi et al., 2024; Goswami et al., 2024). These studies primarily focus on leveraging high-dimensional datasets for pretraining. However, several limitations exist in how they handle such datasets. First, during pretraining, channels are treated independently, capturing only temporal dependencies without modeling inter-channel correlations. Second, despite pretraining on high-dimensional datasets, fine-tuning and evaluation are still performed on low-dimensional ones. Third, some pretraining datasets have unaligned timestamps, complicating their use in evaluation. These limitations highlight the importance of developing high-dimensional time series datasets specifically tailored for benchmarking and evaluation in time series forecasting research. Table 6 in Appendix A provides a more detailed summarization of the differences between Time-HD and existing works.
Multivariate Time Series Forecasting. Previous methods for modeling multivariate time series can be broadly categorized into channel-independent (CI) and channel-dependent (CD) strategies. The CI modeling strategy employs a shared model backbone across all channels and processes each channel independently during the forward pass (Nie et al., 2023; Zeng et al., 2023; Xu et al., 2023; Wang et al., 2024a; Tan et al., 2024; Dai et al., 2024; Lin et al., 2024; Li et al., 2023). This design typically results in lower model capacity but offers greater robustness. In contrast, the CD modeling strategy introduces dedicated modules to capture inter-channel dependencies. Based on the granularity at which these correlations are modeled, CD can be further divided into position-wise(Chen et al., 2023; Wu et al., 2022; 2021; Zhou et al., 2021; 2022; Zhang & Yan, 2023), token-wise(Liu et al., 2024a; Wang et al., 2024c), and cluster-wise (Chen et al., 2024; Qiu et al., 2024b; Li & Anastasiu, 2025) approaches. These methods are expected to provide higher capacity and make better use of cross-channel information. However, due to the constraints of existing low-dimensional time series forecasting benchmarks, most methods have not been validated at scale.
@@ 3 Preliminary Study
Prior studies on existing time series benchmarks, which are often of low dimensionality, do not consistently show a clear advantage of CD over CI models (Zhao & Shen, 2024; Nie et al., 2023; Zeng et al., 2023; Dai et al., 2024). This observation might lead to skepticism regarding the practical benefits of explicitly modeling inter-channel correlations. However, we hypothesize that the limited dimensionality of these benchmarks inherently restricts the potential gains from CD approaches. This preliminary study aims to theoretically and empirically investigate the impact of data dimensionality on the forecasting performance achievable by leveraging channel correlations. We begin with a theoretical analysis of risk reduction.
@@ 3.1 Theoretical Analysis of Risk Reduction from CD Models
To formally analyze the benefits of CD, we consider a simplified setting. Assume a bivariate time series ($C=2$) generated by a Vector Autoregression of order 1 (VAR(1)) (Stock & Watson, 2001):
@@
For the time series described in Eq. equation 1, the Bayes risks of CI and CD models under squared-error loss satisfy: $R_{\text{CI}}-R_{\text{CD}}\geq 0$. This inequality is strict if and only if $a_{12}\neq 0$ and $a_{21}\neq 0$, and the conditional variances $\mathrm{Var}(z_{t}^{(1)}\mid z_{t}^{(2)})$ (representing the variance of $z_{t}^{(1)}$ that remains after $z_{t}^{(2)}$ is known) and $\mathrm{Var}(z_{t}^{(2)}\mid z_{t}^{(1)})$ are positive.
We provide the proof of this theorem in Appendix G. Theorem 1 establishes that CD modeling is at least as good as CI modeling. A strict advantage for CD arises if: (1) there is genuine predictive information flowing between channels (non-zero $a_{12}$ or $a_{21}$ means one channel directly influences the future of the other), and (2) each channel contains some unique information not present in the other (positive conditional variances). If channels are truly independent in their evolution ($a_{12}=a_{21}=0$) or one is a deterministic function of the other, then $R_{\text{CI}}=R_{\text{CD}}$.
To investigate the impact of increasing dimensionality, we now consider a general $P$-channel VAR(1) process: $z_{t+1}=\mathbf{A}z_{t}+\varepsilon_{t+1}$, where $z_{t}\in\mathbb{R}^{P}$. We focus on forecasting a single target channel, say $Y=z_{t+1}^{(1)}$, which is given by $z_{t+1}^{(1)}=\sum_{j=1}^{P}a_{1j}z_{t}^{(j)}+\varepsilon_{t+1}^{(1)}$, with $\mathrm{Var}(\varepsilon_{t+1}^{(1)})=\sigma_{11}$. Let $R_{p}$ denote the Bayes risk (minimum MSE) for forecasting $Y=z_{t+1}^{(1)}$ using the information from the first $p$ channels at time $t$, i.e., $\{z_{t}^{(1)},\dots,z_{t}^{(p)}\}$.
@@
The proof is provided in Appendix H. Theorem 2 formally shows that incorporating additional informative and non-redundant channels progressively reduces the forecasting risk for a target variable. This underscores the theoretical benefit of using more relevant dimensions (channels), supporting our hypothesis that higher-dimensional data provides greater opportunity for CD models to outperform CI approaches.
@@ 3.2 Empirical Analysis
Data Generation. We generate synthetic multivariate time series using a $C$-dimensional Vector Autoregression process in Equation 1, where the coefficient matrix ${\bf A}\in\mathbb{R}^{C\times C}$ encodes the dependency structure of channels. We focus on two distinct structures for $\bf A$}: (1) Independent: Channels evolve independently of each other i.e., $\bf A$} is a diagonal matrix. For this setting, we use $C=100$ channels. (2) Anti-Self: Channels depend significantly on other channels but not on their own immediate past, i.e., $\bf A$} has zero diagonal entries and non-zero off-diagonal entries. To examine the impact of dimensionality, we generate Anti-Self datasets with $C\in\{100,250,2000\}$. More experimental details are provided in Appendix K.
@@
Results and Discussion. Table 2 shows that CI models perform better in the Independent setting without inter-channel dependencies, while CD models achieve markedly lower MSE in Anti-Self settings, with larger gains as the number of interacting channels grows. These results align with our theory that CD modeling becomes increasingly effective when significant inter-channel dependencies exist, especially in higher-dimensional systems. We further confirm this on real-world data (Appendix N). These findings underscore two critical needs: (1) novel CD models that efficiently and effectively navigate the intricacies of high-dimensional time series, and (2) dedicated high-dimensional forecasting benchmarks with complex channel dependencies for rigorous evaluation. Accordingly, in the subsequent sections, we address these needs by first introducing our U-Cast framework specifically tailored for HDTSF and then presenting Time-HD, a comprehensive benchmark designed to support and galvanize further research in this evolving domain.